OpenStudy (anonymous):

Solve (x+1)^2-28=0, where x is a real number. pleeeaaasse help!!!

OpenStudy (lgbasallote):

expand (x+1)^2 first \[\implies x^2 + 2x + 1 - 28 = 0\] \[\implies x^2 + 2x - 27 = 0\] does that help?

OpenStudy (anonymous):

ohhhh! YES thank you!!!

OpenStudy (lgbasallote):

welcome ^_^

OpenStudy (anonymous):

hold up, what do i do next??

OpenStudy (lgbasallote):

quadratic formula :D

OpenStudy (lgbasallote):

\[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] you know how to use it right?

OpenStudy (anonymous):

ohh got it! thank you!

OpenStudy (lgbasallote):

welcome ^_^